. An interaction can be best characterized as:

a) When two IVs affect one another

b) When the effect of an IV on a DV is conditional upon the level of another IV

c) When the effect of an IV on a DV is unconditional upon the level of another IV

d) The average effect of an IV on the DV, ignoring the other IV

e) When the DV behaves differently for different IVs

4. The SSinteraction can be found by:

  1. SStotal - SSbetween
  2. SStotal – (SScol + SSrow)
  3. SSbetween – (SSwithin + SSrow)
  4. SSwithin – (SScol + SSrow)
  5. SSbetween – (SScol + SSrow)

5. A 3 x 4 x 2 Factorial Design with 5 people per cell would have how many total people?

  1. 20
  2. 60
  3. 80
  4. 120
  5. 240
Answer #1

An interaction is best described as when the effect of one IV on a DV is conditional upon the level of another IV

Correct option :B)

4. SStotal = SScol + SSrow + SSinteraction
SSint = SStot - SScol - SSrow

Correct option: B)

5. In a 3x4x2 factorial design, there are a total of 3*4*2 = 24 conditions. Each condition has 5 people.
Hence, Total Size = 24*5 = 120

Correct option: D) 120